# Observation of the quantum equivalence principle for matter-waves

### v3

@@ Abstract
Einstein’s general theory of relativity is based on the principle of equivalence — in essence, dating back to Galileo — which asserts that, locally, the effect of a gravitational field is equivalent to that of an accelerating reference frame, so that the local gravitational field is eliminated in a freely falling frame. Einstein’s theory enables this principle to extend to a global description of relativistic space-time, at the expense of allowing space-time to become curved, realising a consistent frame-independent description of nature at the classical level. Einstein’s theory has been confirmed to great accuracy for astrophysical bodies. However, in the quantum domain the equivalence principle has been predicted to take a unique form involving a gauge phase that is observable if the wavefunction — fundamental to quantum descriptions — allows an object to interfere with itself after being simultaneously at rest in two differently accelerating frames, one being the laboratory (Newtonian) frame and the other in the freely-falling (Einsteinian) frame. To measure this gauge phase we realise a novel cold-atom interferometer in which one wave packet stays static in the laboratory frame while the other is in free-fall. We follow the relative-phase evolution of the wave packets in the two frames, confirming the equivalence principle in the quantum domain. Our observation is yet another fundamental test of the interface between quantum theory and gravity. The new interferometer also opens the door for further probing of the latter interface, as well as to searches for new physics.
I Introduction
@@
Ever since Galileo and Newton, it was clear that the phenomenon of free-fall in a gravitational field is a paradigm of physics, and must be fully understood as part of any description of nature. Einstein went a step further and used the geometry of space-time to describe gravity and free-fall. In addition, an assertion of paramount importance in physics is that to be consistent one should be able to describe nature from different frames of reference, where once more the freely-falling frame is considered fundamental. This led to the formulation of the equivalence principle (EP), which again deals with gravity [1, 2, 3]. The rise of quantum mechanics has not changed any of these convictions, and it remains to be seen how quantum theory adheres to them. The connection between gravity (or general relativity) and quantum theory, the two pillars of modern physics, remains one of the most important open questions in physics.
@@
Here we present a direct observation of the quantum phase of a free-falling object, confirming the quantum equivalence principle (QEP) for matter-waves. Measuring gravity with quantum interferometers has been previously achieved with neutrons in the celebrated Colella-Overhauser-Werner experiment [4, 5] (as well as by Bonse and Wroblewski for a noninertial frame [6]), and with atoms [7, 8, 9, 10, 11], but the explicit phase of a free-falling wave packet (WP) has not been directly measured. This requires a unique interferometer in which one trajectory is stationary in the lab frame (relative to Earth) and acts as a reference, while the other is in free fall.
The most common definition of the EP, as Einstein originally thought of it, is: A state of rest in a homogeneous gravitational field is physically indistinguishable from a state of uniform acceleration in a gravity-free space [12]. Let us note that numerous possible extensions of the equivalence principle to the quantum domain (hereafter, QEP) have been proposed [13, 14, 15, 16, 17, 18, 19, 20] and here we focus on the most straightforward aspect, which is the quantum phase connecting between the freely-falling frame and the lab frame [21], termed by some the Einsteinian and Newtonian frames, respectively [22]. This phase has a characteristic dependence of $t^{3}$ on the free-fall time $t$[3, 21, 22]. As this phase dependence is based on the assumption that the inertial mass ($m_{i}$) is equal to the gravitational mass ($m_{g}$), as suggested by Galileo and as confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [23, 24, 25], we name our unique interferometer the Quantum Galileo Interferometer, or QGI [26]. With the latter assertion, the EP is manifested in a free-falling body under gravity as viewed from accelerated and non-accelerated frames [27], or as Einstein put it: When a person is free-falling they do not feel their weight.
The prediction that an accelerated object, and specifically a free-falling object acquires a $t^{3}$-phase, has been made quite some time ago by Darwin, Kennard and others [28, 29, 30, 31, 32, 33, 34, 35, 36]. While a cubic phase dependence on the interferometer time may also appear in a few standard-practice interferometers [37, 38, 39, 40, 41], observing this free-fall phase in the QGI configuration sampling these two fundamental frames, constitutes a direct test and confirmation of the phase of a free-falling object, as well as a fundamental aspect of the QEP for matter waves [3, 18, 21].
The transformation between the wavefunction $\psi_{N}$ in a Newtonian (laboratory) frame of reference, where the acceleration of an object is due to gravity, $\ddot{z}=-g$, and the wavefunction $\psi_{E}$ in an Einsteinian frame that falls in gravity (with acceleration $a=-g$), and when at $t=0$ both have the same velocity, involves a phase (e.g., Eq. 5 in Ref. [22]) which may also be referred to as a gauge phase $\phi_{\rm gauge}$[21] [see supplementary material (SM)]
According to the equivalence principle, $\psi_{E}$ is the wave function of a free particle that feels no forces in the falling (Einsteinian) frame (as in addition to the force $m_{g}g$ it feels a force $m_{i}a=-m_{g}g$). It is thus exactly equivalent to a particle in the lab (Newtonian) frame if gravity would be cancelled. However, Eq. (1) implies that such a particle accumulates a phase $\phi_{\rm gauge}$ in the lab frame in the presence of gravity. Measuring this fundamental phase thus confirms an outcome of the equivalence principle in the quantum domain.
Our experiment intended to measure this fundamental phase — by way of interference with a reference WP — follows a Gedanken (thought) experiment where at time $t=t_{0}$ the WP is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied force that is exactly opposite to gravity, while in the other path it is in free-fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the WP in the freely falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter referred to as the closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two WPs so that the two paths overlap in position and momentum. Fig. 1(A) describes this Gedanken experiment, where the WP that is at rest in the lab (Newtonian) frame is termed the “reference WP” while the WP that is free-falling in the lab (Newtonian) frame is termed the “ballistic WP” [42].
@@
For an intuitive understanding of the QGI in the Newtonian frame, we fix the reference WP center at vertical position $z=0$ so that it does not accumulate any propagation phase (beyond the free evolution of the WP), and calculate the phase accumulated by the ballistic WP. The latter accumulates no propagation phase in the freely-falling (Einsteinian) frame, and we can use the gauge phase to know its phase in the Newtonian frame. We calculate the phase in the ballistic path in two parts: First, a WP starting at $t=0$ in the middle of the interferometer and going forward in time, and then starting from the same point and going backward in time, with these paths ending at $t=\pm T$ and $z=0$ after a free-fall duration of $T$ [Fig. 1(A)]. It follows that the gauge phase at the center of the WP for the two parts, and the phase difference between the two end points read
$\phi_{\rm gauge}(\pm T)=\mp\frac{m}{6\hbar}g^{2}T^{3}\,{\rm;}\,\,\phi_{\rm ballistic}=\phi_{\rm gauge}(T)-\phi_{\rm gauge}(-T)=-2\cdot\frac{m}{6\hbar}g^{2}T^{3}\,,$ (2)
where the linear phase contribution in Eq. (1) is canceled by the magnetic-gradient pulses (SM). In the SM we arrive at the same result also by starting from the Galilean transformation. Consequently, $\phi_{\rm ballistic}$ is the relative phase observable in the experiment.
In fact, the QGI also demonstrates a QEP between the Newtonian and Einsteinian frames, for a quantum system in a state of spatial superposition. This is due to the complete symmetry of the QGI. An observer sitting in the Einsteinian or Newtonian frames will see exactly the same phenomenon, with the roles of the reference WP and the ballistic WP inverted. This symmetry becomes strikingly clear when one computes the actions in both frames (SM). Assuming that the inertial and gravitational masses are equal and that the acceleration we induce by an external force to form the reference trajectory is exactly opposite to the gravitational acceleration, the action of the ballistic path in the Newtonian frame and that of the reference path in the Einsteinian frame are the same, equal to $-\frac{m}{3}g^{2}T^{3}$[43].
Finally, the QGI is able to measure the QEP for matter waves, as it utilizes no electromagnetic waves resonant with some atomic transition. In contrast, in interferometers based on light pulses, one must take into account the electromagnetic fields, namely, introduce transformations of Maxwell equations, so that a working interferometer may be described in both frames. Consequently, in such interferometers, what is in fact verified is the EP for electromagnetic waves [44, 45] (see SM). In the QGI, no such additional theoretical structure needs to be invoked, so that only transformations of de-Broglie waves are required to understand the QGI in all frames.
To conclude, it may then be understood that the QGI is able to observe the QEP for matter waves due to two main reasons: First, it is a configuration in which one WP is at rest in an inertial frame while the other is at rest in a freely-falling frame, and second, by only utilizing the theory of de-Broglie waves it achieves the complete symmetry required for observing the QEP for matter waves.
Figure 2: Population in outport 1 vs. the free-fall duration of the ballistic path $\bm{2T}$.
(A) The relative spin population in outport 1 (state $|0\rangle$) oscillates with a distinct chirp as a function of interferometer time. The blue points represent the experimental data, with error bars derived from the standard error of the mean (SEM). The red line is a fit to the data of the form $P=P_{\text{mean}}(2T)+\frac{1}{2}V(2T)\cos\left[\phi(2T)\right]$, where $V$ is the visibility of the oscillations, $\phi$ is the phase of the oscillations and $P_{\text{mean}}$ is the mean value of the population. To obtain the fit, we first identify the upper and lower envelopes and fit them to a polynomial to get $P_{\text{mean}}$ and $V(2T)$. We then fit the data to the model where the phase $\phi(2T)$ is a third-order polynomial. During the fitting we exclude the first and last oscillation as they suffer from larger uncertainty in the values of the envelope. The data shows 13 complete oscillations, with low phase noise, resulting in an average population error (SEM) of $1.8\pm 1.5$ [%] and good visibility, starting at $V=80\,\%$ and decreasing to $V=20\,\%$ at $2T=2000\,\mu\mathrm{s}$. The decrease in visibility is mainly due to a varying overlap between the WPs due to differences in shape evolution under the influence of curved gradients as well as imperfection in the path recombination. The figure includes $633$ experimental cycles, each $30$ s, so the graph was taken over a period of $5.3$ hours. The short-term and long-term stability of the interferometer are good due to the high stability of the chip trap, and current source. (B) Atomic distribution (heat map of optical density) adapted from CCD images (on-resonance absorption imaging), of the atoms in the two interferometer outports [see Fig. 1(B)], showing two oscillations. The top cloud is outport 1 (state $|0\rangle$), and the bottom cloud is outport 2 (state $|1\rangle$, which is magnetically sensitive). The cloud in outport 2 is focused by the final splitting pulse, increasing its optical density, and making it appear slightly denser even when the two clouds have the same population. The vertical scale of the image is 1.1 mm.
II Experiment
Typical atom interferometers are free-space interferometers, whereby during most of their motion, the WPs do not experience any potential applied by the experimental setup [46]. Some interferometers do have a potential applied to the WPs during their evolution, and these may be termed, guided interferometers [47, 48]. In our case, both methods are used simultaneously, one for each of the WPs.
In Fig. 1(B) we present the experimental setting in which the QGI, utilizing 87Rb atoms, is conducted about $113\,\mu$m below an atom chip [49], where the current-carrying wires give rise to the magnetic field gradients which are responsible for the Stern-Gerlach (SG) type forces applied to the WPs. The QGI is a novel type of SG interferometer (SGI), leveraging a decade of SGI experiments in which the necessary techniques were developed [50], and consequently, is now capable of demonstrating enhanced stability, larger phase accumulation, and especially the flexibility required to enable one WP to take a ballistic trajectory while the other is stationary.
@@ II Experiment
During the free-fall time, the ballistic WP is in state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, so that to first order in the Zeeman interaction it is not affected by the magnetic gradients. The reference WP is in state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and is held stationary by a “holding magnetic gradient pulse” producing a magnetic force counteracting gravity. Both states belong to the atomic ground state, whereby $F$ is the total angular momentum of an atomic state and $m_{F}$ is its projection on the quantization axis.
@@
The splitting of the WP is not abrupt and requires a finite kick time $T_{kick}$ in which the reference WP is temporarily in state $|0\rangle$ and the ballistic WP is temporarily in state $|1\rangle$ allowing it to be accelerated with an average acceleration $a_{kick}=v_{0}/T_{kick}$. In addition, a finite delay time $T_{d}$ is needed between the kick pulse and the holding pulse of duration $T_{h}=2T-2T_{d}$, where 2T represents the free-fall time of the ballistic WP. As explained in Methods, the ballistic WP has a free-fall time of $2T$, while the reference WP is in free fall during the time $T_{kick}+T_{d}$ before and after the holding time. Using $\phi_{gauge}$ these latter parts of the reference WP trajectory involve a free-fall phase in the Newtonian (lab) frame of
@@
(A) The phase difference between the two interferometer arms as a function of the free-fall duration of the upper arm. The red line represents the experimental data, where the phase is extracted from Fig. 2. To estimate the uncertainty bounds, we repeat the experiment with an increase (decrease) of the magnetic-pulse current $I_{kick}$ by $0.5\,\%$ (which is our estimated experimental precision) and use the result to plot the upper (lower) uncertainty bounds. The green points are the result of a numerical simulation based on tools we developed for calculating WP evolution [54], which include the effects of shape and rotation of the WP. The blue dashed line is the result of the analytical approximation of the experiment given in Eq. 4. The analytical model assumes a homogeneous magnetic gradient, and a set of square pulses of current so that the acceleration of each arm is a piecewise constant function. The black dashed line depicts the phase of the Gedanken experiment. The inset shows the derivative of the phase with respect to $2T$. To obtain the derivative of the phase for the numeric simulation, we fit a third-order polynomial to the set of points in the main figure and take its derivative. The derivative of the analytical approximation and the Gedanken experiment are calculated by direct differentiation. The yellow dashed lines represent 5% deviations from the $T^{3}$ phase accumulation model, by modifying the analytical prediction to a power law of $T^{3\pm 0.15}$. This is meant to show that the data is not consistent with such deviations from the cubic phase. (B) The residuals between the phase of the data and the numerical simulation, where the latter is a “blind” simulation with no free parameters or fine tuning. A difference of 2 rad over 13 oscillations spanning a phase of $\approx 80$ rad, constitutes a deviation of about 2.5%. The error bars represent the statistical noise of the phase in the experiment, which shows a relative noise of 1.3% of the phase. The red band represents the systematic uncertainties [which also appear in (A)]. The blue line shows the result of the numeric simulation after fine-tuning of the currents and the magnetic field along the z-axis. The kick current is increased by 0.15% , the idle current is changed from 0.55 mA to 0.5 mA, and $B_{z}$ from 0.33 G to 0.35 G, all within the experimental uncertainties.
@@
Beyond the achievement of two very different WP trajectories in space, a significant experimental challenge lies in the fact that the two trajectories are very much different in momentum, and recombining the WPs in order to observe an interference pattern requires a good overlap in both position and momentum (namely, zeroing any difference, or in other words, time-reversing the splitting process). The time-irreversibility of the splitting process, whether it be of technical or fundamental origin, has been termed by Scully, Englert and Schwinger the “Humpty-Dumpty effect” and combating it requires meticulous control over the trajectories [51].
@@
More so, for a good overlap the size and shape of the two WPs have to be quite similar, and as they are exposed to different magnetic gradients having different curvature, lensing effects make such an overlap hard to achieve. Finally, also rotation of the WPs in 3D may diminish the eventual overlap, and consequently the contrast (visibility) of the interference pattern [52].
@@
Whereas recombination in atom interferometers based on laser pulses, even when the momentum splitting is large [53], enjoys the naturally given quantum accuracy of the photon momentum, in an SGI the momentum transfer is classical in nature. This requires a high level of control over the magnetic fields as well as a detailed simulation of the WP evolution when exposed to such fields and gravity [54]. The QGI indeed achieves good control over position, momentum, acceleration, as well as WP shape and spin.
III Results

### v4

# Observation of quantum free fall and the consistency with the equivalence principle

@@ Abstract
The unification of quantum theory and the general theory of relativity – describing gravity, is one of the most important challenges in science. Einstein’s general theory of relativity is based on the principle of equivalence, and has been confirmed to great accuracy for large bodies. However, in the quantum domain the equivalence principle has been predicted to take a unique form involving a gauge phase, equal to the quantum phase of a free-falling object. To measure this phase, we realize a novel cold-atom interferometer in which one wave-packet stays static in the laboratory frame while the other is in free fall. The observed relative-phase of the wave-packets confirms the predicted phase of a free-falling object, and shows that in our low energy regime, the equivalence principle may be applied to the quantum domain. Our observation constitutes a fundamental test of the interface between quantum theory and gravity. The new interferometer also opens the door for further probing of the latter interface, as well as to searches for new physics.
1 Introduction
@@
The two pillars of twentieth-century physics—quantum mechanics (QM) and general relativity (GR)—have long stood side by side, resisting all attempts to bring them convincingly together. Related to the latter, ever since Galileo and Newton it was clear that the phenomenon of free fall in a gravitational field is a paradigm of physics and must be fully understood as part of any description of nature. Einstein went a step further and used the geometry of space-time to describe gravity and free fall. An additional outcome of GR is that one should be able to describe nature from different frames of reference. Rooted in Galileo’s insight and elevated by Einstein to a foundational axiom, the principle of equivalence has guided our understanding of gravitation for over a century [1, 2, 3, 4]. The rise of quantum mechanics has not changed any of these convictions. The connection between these two pillars of modern physics, remains one of the most important open questions in physics, and it is thus of paramount importance to understand the phenomenon of free fall and the equivalence principle (EP) also in the quantum domain.
The phase of free fall is predicted in a purely quantum manner to have a dependence $\frac{m}{6}g^{2}T^{3}/\hbar+gmzT$ on the free-fall time $T$, where $m$ is the mass of the object, $g$ is the gravitational acceleration relative to the surface of Earth, and $z$ in the spatial coordinate in the direction of gravity. This prediction follows the calculated phase accumulated by an object accelerating in a linear potential, and has been made starting from almost one hundred years ago by Darwin, Kennard and others [5, 6, 7, 8, 9, 10, 11, 12, 13]. Alternatively, one may derive this phase in a completely different manner, specifically, utilizing the EP and a Galilean transformation. The identity of the outcome of the two calculations may be seen as what allows QM and the EP to coexist. Indeed, numerous authors have suggested that this phase has implications regarding the EP in the quantum domain [8, 3, 14, 15, 16]. Surprisingly, this crucial phase has never been measured.
@@
Matter-wave quantum interferometers are high-precision tools capable of measuring global and relative accelerations. Such interferometry in the presence of the gravitational field of Earth has been previously achieved with neutrons in the celebrated Colella-Overhauser-Werner (COW) experiment [17, 18] (as well as by Bonse and Wroblewski for a noninertial frame [19]), and with atoms [20, 21, 22], where the gravitational acceleration gg[23, 24, 25, 26] and the gravitational constant GG[27, 28] have been measured, but as noted, the explicit phase of a free-falling wave packet (WP) with respect to Earth has not been directly measured.
Here we report the measurement of the quantum phase of an object in free fall (in the sense that the experimentalist does not apply any forces to it), as observed in a frame stationary with respect to Earth. A direct measurement of this phase requires a unique interferometer employing a superposition of one WP in free fall, and another WP stationary in Earth’s frame. We have named this interferometer the Quantum Galileo Interferometer, or QGI [29].
Fig. 1(A) describes such a Gedanken (thought) experiment, where the WP that is at rest in the laboratory (Earth) frame is termed the “reference WP” while the WP that is free falling in this frame is termed the “ballistic WP” [30].
@@
While a cubic phase dependence on the interferometer time may also appear in a few situations that are not related to free fall [31, 32, 33, 34, 35], observing this free-fall phase in the QGI configuration constitutes a direct test and confirmation of the phase of a free-falling object, and as noted, of the predicted phase when applying the EP to a quantum WP in free fall. Specifically, none of the previous experiments measuring a $T^{3}$ behavior, have had in the two interferometer arms the accelerations of 0 and $g$ with respect to Earth, and consequently, none measured the predicted prefactor of $\frac{m}{6}g^{2}/\hbar$. As explained in the following, this requires a hybrid interferometer to be realized, a very different configuration from the latter interferometers.
To briefly show how the same phase is predicted by the EP, we will assume that the EP holds in the quantum domain, and calculate the observable consequence of such an assumption. To begin with, throughout this work we assume the equality of the inertial mass ($m_{i}$) and the gravitational mass ($m_{g}$), such that $m_{i}=m_{g}=m$, as suggested by Galileo and confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [36, 37, 38, 39]. We of course also assume the universality of free fall to hold, so that also a $^{87}Rb$ atom in free fall accelerates towards Earth with acceleration $g$.
We now view the experiment from two frames of reference, one in free fall and one static relative to Earth. These two frames may be referred to as the Einsteinian and the Newtonian frames, respectively [15]. The transformation between the wavefunction $\psi_{N}$ in a Newtonian (laboratory) frame of reference, where the acceleration of a free object is denoted as $\ddot{z}=-g$, and the wavefunction $\psi_{E}$ of the same object in an Einsteinian frame that free-falls with the same acceleration g-g, and when at $t=0$ both have the same velocity, involves a transformation phase which may also be referred to as a gauge phase $\phi_{\rm gauge}$[16]. This gauge phase may be calculated in ways that do not rely on the EP, for example, by using Galilean transformations [see Supplementary Material (SM)].
If one then assumes that the inertial mass is equal to the gravitational mass, this transformation phase takes the form (e.g., [15, 16]):
According to the EP, $\psi_{E}$ is the wavefunction of a free particle that feels no forces in the free-falling (Einsteinian) frame [40]. Hence, if one assumes that the EP holds, this WP does not accumulate any phase (beyond free evolution). However, Eq. (1) implies that such a wavefunction accumulates a phase $\phi_{\rm gauge}$ in the laboratory (Newtonian) frame, simply due to the transformation between frames.
In conclusion, when comparing a quantum state in two different descriptions—one in which the gravitational field is treated as an external force, and another adopting the freely falling frame—a unique phase factor emerges, $mg^{2}t^{3}/6\hbar$[8, 3, 14, 15, 16]. This contribution, first identified nearly a century ago as the phase of a freely-falling object, has since attracted considerable theoretical scrutiny [5, 6, 7, 8, 9, 10, 11, 12, 13]. However, while this phase was typically calculated as the phase of an accelerating particle, without regard for the problem of the interplay between QM and GR, remarkably, as highlighted by the numerous works [8, 3, 14, 15, 16], without this phase QM and the EP do not coexist. Specifically, by using the EP, as we did above, one predicts the same phase. In the case that the quantum calculation of an accelerating object under gravity, as was calculated by the numerous authors, was not complete, or in case the application of the EP to the quantum domain, as we have done, is lacking, it is of paramount importance to experimentally confirm this phase. We emphasize that in this work we will not be excluding all possible models for which the EP does not hold, as such models may also predict the same phase. Assuming the EP holds and observing no contradiction, we are able to show that the EP may be applied to a quantum WP. Hence, beyond the measurement of the phase of a free-falling quantum object, the QGI enables the observation of the phase enabling the coexistence of QM and the EP.
2 Experiment
Our experiment intended to measure this phase — by way of interference with a reference WP — follows the Gedanken experiment presented in Fig. 1(A) where at time $t=t_{0}$ the WP is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied levitating force that is exactly opposite to gravity (hereafter levitation condition), while in the other path it is in free fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the WP in the freely-falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two WPs so that the two paths overlap in position and momentum.
@@ 2 Experiment
For an intuitive understanding of the QGI phase in the Newtonian frame, we fix the reference WP center at vertical position $z=0$ so that it does not accumulate any propagation phase (beyond the free evolution of the WP), and calculate the phase accumulated by the ballistic WP. As noted, the latter accumulates no propagation phase in the freely-falling (Einsteinian) frame, so that the phase at the center of the ballistic WP in the Newtonian frame is nothing but the gauge phase. The phase of the ballistic path is the difference between the phase of the WP center at the two path endpoints, which is equal to the gauge phase at these points. Taking the middle of the path to be at time $t=0$, these endpoints are at $t=\pm T$, where the WP is centered at $z=0$ after a free-fall duration $T$ with respect to $t=0$ (Fig. 1(A)). The gauge phase at the endpoints and the phase of the ballistic path then read
$\phi_{\rm gauge}(\pm T)=\mp\frac{m}{6\hbar}g^{2}T^{3}\,{\rm;}\,\,\phi_{\rm ballistic}=\phi_{\rm gauge}(T)-\phi_{\rm gauge}(-T)=-2\cdot\frac{m}{6\hbar}g^{2}T^{3}\,.$ (2)
As $\phi_{\text{reference}}=0$, it follows that $\phi_{\text{ballistic}}$ is the relative phase observed in the experiment [41]. This phase is of course equal to the action of the ballistic path in the Newtonian frame, or equivalently of the reference path in the Einsteinian frame, as we show in the SM. In the SM, we also show that it can be calculated using the Galilean transformation. Note also that the linear term in Eq. (1) vanishes when the endpoints are at $z=0$ and even if the levitation condition is not satisfied it is cancelled by the magnetic gradient pulses, which are necessary for the closing condition (SM).
Beyond the above description of the fundamental uniqueness of the QGI, in which the measured phase is equal to the free-fall phase, its technical distinctiveness should also be emphasized. While the Ramsey-Bordé interferometer is mostly characterized by the constant spatial separation between the two interferometer paths [42], and the Kasevich-Chu interferometer is mostly characterized by the piece-wise constant velocity separation (difference) between the two paths [43], the QGI enables the control of all degrees of freedom, obtaining spatial, velocity and acceleration differences, almost at will. This allowed us to have one WP at rest and the other free falling. Furthermore, typical atom interferometers are free-space interferometers, whereby during most of their motion, the WPs do not experience any potential applied by the experimental setup [44]. Some interferometers do have a potential applied to both WPs during their evolution, and these may be termed, guided interferometers [45, 46]. In our case, both methods are used simultaneously, one for each of the WPs. The QGI is thus a hybrid interferometer.
[…]
@@
During the free-fall time, the ballistic WP is in state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, so that to first order in the Zeeman interaction it is not affected by the magnetic gradients. The reference WP is in state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and is held stationary by a “holding magnetic gradient pulse” producing a magnetic acceleration $a$ in the opposite direction to gravity. Both states belong to the atomic ground state, whereby $F$ is the total angular momentum of an atomic state and $m_{F}$ is its projection on the quantization axis. The acceleration $a$ is independently tuned by an experimental procedure (see Methods and SM) to satisfy the levitation condition for which the reference WP is stationary in the laboratory frame (i.e., relative to Earth). In case of small experimental errors, one can consider the general case that includes deviations from the levitation condition, for which a simple calculation of the actions along the two interferometer trajectories when the closing condition is satisfied results in a measured QGI phase
$\Delta\phi=\frac{ma(a-2g)}{3\hbar}T^{3}\,$, (3)
such that the outcome of the QGI depends on both the applied external acceleration on the reference arm and the free-fall acceleration, and becomes equal to the result of Eq. (2) in the case $a=g$, with a quadratic deviation $m(a-g)^{2}T^{3}/3\hbar$ if $a\neq g$. While the measurement of $g$ is not the aim of our experiment, we should note that the sensitivity with which a $T^{3}$ SGI, such as ours, can measure $g$ was recently estimated [49].
@@
The momentum splitting of the WPs in the experiment requires a finite kick duration $T_{kick}$ in which the reference WP is temporarily in state $|0\rangle$ and the ballistic WP is temporarily in state $|1\rangle$ allowing it to be accelerated with an average acceleration $a_{kick}=v_{0}/T_{kick}$. Furthermore, in order to perform the internal-state flip before or after the kick, a finite delay time $T_{d}$ is needed between the kick pulse and the holding pulse of duration $T_{h}=2T-2T_{d}$, where 2T represents the free-fall time of the ballistic WP. As explained in Methods, the ballistic WP has a free-fall time of $2T$, while the reference WP is in free fall during the time $T_{kick}+T_{d}$ before and after the holding time. Using $\phi_{gauge}$ these latter parts of the reference WP trajectory involve a free-fall phase in the laboratory (Newtonian) frame of
@@
Beyond the achievement of two very different WP trajectories in space, a significant experimental challenge lies in the fact that the two trajectories are very much different in momentum, and recombining the WPs in order to observe an interference pattern requires a good overlap in both position and momentum (namely, zeroing any difference, or in other words, time-reversing the splitting process). The time-irreversibility of the splitting process, whether it be of technical or fundamental origin, has been termed by Scully, Englert and Schwinger the “Humpty-Dumpty effect” and combating it requires precise control over the trajectories [50].
@@
More so, for a good overlap the size and shape of the two WPs have to be quite similar, and as they are exposed to different magnetic gradients having a different curvature, lensing effects make such an overlap hard to achieve. Finally, also rotation of the WPs in 3D may diminish the eventual overlap, and consequently the contrast (visibility) of the interference pattern [51].
@@
Whereas recombination in atom interferometers based on laser pulses, even when the momentum splitting is large [52], enjoys the naturally given quantum accuracy of the photon momentum, in an SGI the momentum transfer is classical in nature. This requires a high level of control over the magnetic fields as well as a detailed simulation of the WP evolution when exposed to such fields and gravity [53].
3 Results